# Rotatable Antenna-Enhanced Cell-Free Communication

### v1

@@ Abstract
Rotatable antenna (RA) is a promising technology that can exploit new spatial degrees-of-freedom (DoFs) by flexibly adjusting the three-dimensional (3D) boresight direction of antennas. In this letter, we investigate an RA-enhanced cell-free system for downlink transmission, where multiple RA-equipped access points (APs) cooperatively serve multiple single-antenna users over the same time-frequency resource. Specifically, we aim to maximize the sum rate of all users by jointly optimizing the AP-user associations and the RA boresight directions. Accordingly, we propose a two-stage strategy to solve the AP-user association problem, and then employ fractional programming (FP) and successive convex approximation (SCA) techniques to optimize the RA boresight directions. Numerical results demonstrate that the proposed RA-enhanced cell-free system significantly outperforms various benchmark schemes.
@@ I Introduction
With the rapid development of global information and communication technology (ICT), the denser deployment of access points (APs) has become a prevailing trend in communication systems to meet the increasing demands for lower latency, higher reliability, and improved spectral efficiency. A promising paradigm is the cell-free multiple-input multiple-output (MIMO) system [1]. In this system, numerous APs, distributed throughout the service area and connected to a central processing unit (CPU) via fronthaul, collaboratively serve all users without cell boundaries [2]. However, conventional cell-free MIMO systems typically employ isotropic and fixed antennas, lacking the capability to dynamically distribute antenna resources in the spatial domain. This inherent limitation consequently restricts the communication performance.
To overcome this limitation, rotatable antenna (RA) [5, 6, 7, 8, 9] has recently emerged as a promising technology to exploit new spatial degrees-of-freedom (DoFs) by dynamically adjusting the antennas’ three-dimensional (3D) boresight directions. As a simplified yet appealing variant of the six-dimensional movable antenna (6DMA) [3, 4] paradigm leveraging only the rotation flexibility (without changing antenna positions), RA enables a cost-effective and compact design more suitable for practical deployment [5]. Due to these appealing advantages, recent studies have explored the performance gains enabled by RA in various wireless systems. The authors in [6] and [7] established the fundamental system model and channel characteristics of RA systems, demonstrating their ability to achieve more favorable channel conditions. Building on this, RA has been applied to more challenging scenarios, such as integrated sensing and communication (ISAC)[8] and physical-layer security[9]. Notably, through flexible 3D boresight control, RA can enhance the signal in desired directions, thereby making it well-suited for improving the communication performance of cell-free systems. Nevertheless, the application of RA in such systems remains unexplored, thus motivating further investigation in this work.
@@ II-B Signal Model
As a prerequisite for precoding and subsequent optimization, we assume that the global channel state information (CSI) can be fully acquired by the CPU via the fronthaul[10, 11]. In addition, to achieve coherent superposition and maximize the received signal power, we adopt normalized conjugate beamforming with precoding factor $\phi_{l,k}^{*}\triangleq\frac{h_{l,k}^{*}}{\left|h_{l,k}\right|}$. Thus, the downlink transmitted signal $s_{l}$ at AP $l$ can be written as
@@ III-A AP-User Association Strategy
By exploiting the fact that nearer AP–user pairs suffer less attenuation and are more likely to achieve better performance, we form an approximate substitution to (P1). Specifically, by calculating the distances $d_{l,k}$ of all possible AP-user pairs, we can reformulate the original association problem as minimizing the total AP-user association distance $D_{\text{sum}}=\sum_{l=1}^{L}{\sum_{k=1}^{K}{b_{l,k}d_{l,k}}}$. Accordingly, we have:
@@
The proposed algorithm ensures that each user is served by at least one AP and that no AP remains idle. It should be noted that (P2) is an approximate surrogate of (P1), obtained by omitting the effect of small-scale fading. Although this leads to suboptimal solutions, Algorithm 1 provides an effective AP-user association with low computational complexity, requiring only basic operations such as sorting and comparison. Specifically, the complexity order for solving problem (P2) is $\mathcal{O}\left(KL\right)$.
@@ III-B RA Pointing Vector Optimization
It can be verified that problem (P6) is convex and can be solved by the CVX solver[13]. Note that the optimal value obtained by problem (P6) serves as an upper bound for that of problem (P5) due to the relaxation of the equality constraint  (16b). After obtaining the optimal solutions of $\vec{\mathbf{f}}_{l}$ through Algorithm 2, the pointing vector needs to be recovered as a unit vector, i.e., $\vec{\mathbf{f}}_{l}^{\star}=\frac{\vec{\mathbf{f}}_{l}}{\lVert\vec{\mathbf{f}}_{l}\rVert}$. This normalization only scales $\vec{\mathbf{f}}_{l}$ to unit length without altering its direction.
@@
until The fractional increase of the sum rate in (17a) falls below the threshold $\xi$ or the iteration number $i$ reaches the pre-designed number of iterations II.
@@ IV Simulation Results
Figure 4: Sum rate (bps/Hz) vs. the number of APs where $K=5$.
@@
Figure 5: Average per-user rate (bps/Hz) vs. the number of users where $L=30$.

### v2

@@ Abstract
Rotatable antenna (RA) is a promising technology that can exploit new design dimensions by optimizing the three-dimensional (3D) boresight direction of directional antennas. In this letter, we investigate an RA-enhanced cell-free system for downlink transmission, where multiple RA-equipped access points (APs) cooperatively serve multiple single-antenna users over the same time-frequency resource. Specifically, we aim to maximize the sum rate of all users by jointly optimizing the AP-user associations and the RA boresight directions. Accordingly, we propose a two-stage strategy to solve the AP-user association problem, and then employ fractional programming (FP) and successive convex approximation (SCA) techniques to optimize the RA boresight directions. Numerical results demonstrate that the proposed RA-enhanced cell-free system significantly outperforms various benchmark schemes.
@@ I Introduction
With the rapid development of global information and communication technology (ICT), the denser deployment of access points (APs) has become a prevailing trend in communication systems to meet the increasing demands for lower latency, higher reliability, and improved spectral efficiency. A promising paradigm is the cell-free multiple-input multiple-output (MIMO) system [1], where numerous APs, distributed throughout the service area and connected to a central processing unit (CPU) via fronthaul, collaboratively serve all users and thereby alleviate cell boundaries [2]. However, conventional cell-free MIMO systems typically employ isotropic and fixed antennas, lacking the capability to dynamically distribute antenna resources in the spatial domain. This inherent limitation consequently restricts the communication performance.
To overcome this limitation, rotatable antenna (RA) [5, 6, 7, 8, 9] has recently emerged as a promising technology to exploit new design dimensions by dynamically adjusting the antennas’ three-dimensional (3D) boresight directions. As a simplified yet appealing variant of the six-dimensional movable antenna (6DMA) [3, 4] paradigm leveraging only the rotation flexibility (without changing antenna positions), RA enables a cost-effective and compact design more suitable for practical deployment [5]. Due to these attractive advantages, recent studies have explored the performance gains enabled by RA in various wireless systems. The authors in [6] and [7] established the fundamental system model and channel characteristics of RA systems, demonstrating their ability to achieve more favorable channel conditions. Building on this, RA has been applied to more challenging scenarios, such as integrated sensing and communication (ISAC)[8] and physical-layer security[9]. Notably, through flexible 3D boresight control, RA can enhance the signal in desired directions, thereby making it well-suited for improving the communication performance of cell-free systems. Nevertheless, the application of RA in such systems remains unexplored, thus motivating further investigation in this work.
@@ II-B Signal Model
As a prerequisite for precoding and subsequent optimization, we assume that global channel state information (CSI) can be fully acquired by the CPU via the fronthaul[10, 11]. In addition, to achieve coherent superposition and maximize the received signal power, we adopt normalized conjugate beamforming with precoding factor $\phi_{l,k}^{*}\triangleq\frac{h_{l,k}^{*}}{\left|h_{l,k}\right|}$. Thus, the downlink transmitted signal $s_{l}$ at AP $l$ can be written as
@@ III-A AP-User Association Strategy
By exploiting the fact that pairing nearby APs and users causes less link attenuation and is more likely to achieve high performance, we approximate (P1) for AP-user pairing. Specifically, by calculating the distances $d_{l,k}$ of all possible AP-user pairs and ignoring interference, we reformulate the original association problem as minimizing the total AP-user association distance $D_{\text{sum}}=\sum_{l=1}^{L}{\sum_{k=1}^{K}{b_{l,k}d_{l,k}}}$. Accordingly, we have:
@@
The proposed algorithm ensures that each user is served by at least one AP and that no AP remains idle. It should be noted that (P2) is an approximate surrogate of (P1), obtained by omitting the effects of small-scale fading and interference. Although this leads to a suboptimal solution, Algorithm 1 provides an effective AP-user association with low computational complexity, requiring only basic operations such as sorting and comparison. Specifically, the complexity order for solving problem (P2) is $\mathcal{O}\left(KL\right)$.
@@ III-B RA Pointing Vector Optimization
It can be verified that problem (P6) is convex and can be solved by the CVX solver[13]. Note that the optimal value obtained by problem (P6) serves as an upper bound for that of problem (P5) due to the relaxation of the equality constraint  (16b). After obtaining the optimal solution of $\vec{\mathbf{f}}_{l}$ through Algorithm 2, the pointing vector needs to be recovered as a unit vector, i.e., $\vec{\mathbf{f}}_{l}^{\star}=\frac{\vec{\mathbf{f}}_{l}}{\lVert\vec{\mathbf{f}}_{l}\rVert}$. This normalization only scales $\vec{\mathbf{f}}_{l}$ to unit length without altering its direction.
@@
until The fractional increase of the sum rate in (17a) falls below threshold $\xi$ or iteration number $i$ reaches the pre-designed number of iterations II.
@@ IV Simulation Results
Figure 4: Sum rate (bps/Hz) vs. the number of APs for $K=5$.
@@
Figure 5: Average per-user rate (bps/Hz) vs. the number of users for $L=30$.